# DiffCJK: Conditional Diffusion Model for High-Quality and Wide-coverage CJK Character Generation

### v1

@@ Abstract
Chinese, Japanese, and Korean (CJK), with a vast number of native speakers, has profound influence on society and culture. The typesetting of CJK languages carries a wide range of requirements due to the complexity of their scripts and unique literary traditions. A critical aspect of this typesetting process is that CJK fonts need to provide a set of consistent-looking glyphs for approximately one hundred thousand characters. However, creating such a font is inherently labor-intensive and expensive, which significantly hampers the development of new CJK fonts for typesetting, historical, aesthetic, or artistic purposes.
@@ Introduction
The importance of East Asian languages that use Chinese characters, including Chinese, Japanese and Korean (CJK) languages, could never be overstated. For example, they are used by more than $1.56$ billion speakers (?) which account for more than $25\%$ of global GDP (?). Also, historically, a wide range of scripts have been heavily influenced by Chinese characters, adding more historical and culture value. In the modern era, with the proliferation of printing and information technology, writing scripts need to address the challenges typesetting: for example, the requirements on the visual appearance of characters means that a font must provide well-designed and consistent-looking glyphs across the glyph repertoire.
However, the unique aspect for CJK typesetting is the sheer number of characters that makes such requirement more effort demanding: CJK contains nearly one-hundred thousand unique characters in the latest Unicode 15.1 (?). Furthermore, there is also an emphasis on typefaces categories with different styles [^1] Here “style” is in a general sense rather than its specific meaning in font development as in “style instance”., as they carry practical, cultural, and artistical meaning (???). Consequently, the font industry faces a dilemma: either making one style instance with as wide of coverage as possible, or making fonts with wider style support but with limited coverage. Despite these compromises, font creation remains an laborious and expensive task. Therefore, it remains a critical challenge to efficiently producing fonts that encompass a diverse range of various, multi-style yet legitimate CJK glyphs for the entire set of nearly one hundred thousand characters.
To address these challenge, we propose a diffusion-based method (Figure 1) capable of generating characters in diverse styles conditioned on a single standard reference glyph. The reference glyph only needs to outline the character’s shape, obviating the need for intricate design work. As a diffusion model, it consists of two processes: a diffusion process that gradually adds noise to destroy the data, and a reverse process that generates new samples from scratch by progressively removing noise. A deep neural network is trained to estimate the noise, conditioned on the reference character. Unlike common text-to-image models which must memorize the shapes of characters and thus suffering from limited data of rare characters, our method works uniformly well for all CJK glyphs, from common to rare ones, as long as we have an available reference. Furthermore, by injecting style information in the temporal embedding for diffusion, our method can not only generate characters in various styles but also interpolate between these styles.
@@
The importance of glyphs in typesetting is multifaceted, with style (typefaces) and weight (??) being two crucial aspects. In Latin typesetting, well-recognized styles include the serif (“brown fox”) and sans-serif (“brown fox”), while weight variations are exemplified by bold (“brown fox”) and italics (“brown fox”). CJK typesetting is not an exception, and even encompasses more delicacy regarding this matter: it requires a consistent appearance across a vast number of characters combined with unique cultural practices. Also, CJK typesetting places a significant emphasis on typeface categories, which carry practical, cultural, and artistic significance. (???).
Notably, styles in CJK typesetting, while similar, do not entirely align with Latin typesetting practices. They are better shown by several key styles unique to CJK typography in Figure 3: (1) Ming/Song, a printed form style that has evolved from the regular script over centuries, is the most widely used and is akin the serif style in Latin typography. (2) Gothic typefaces, printed form styles that convey a sense of modernity and are akin to sans serif styles in Latin typography. (3) Regular script, a printed form style that mimics handwritten forms, is primarily used for educational purposes and also serves a role similar to italics in Latin typography, although italics per se do not exist in CJK typesetting. (4) Calligraphy form, such as semi-cursive and cursive script, that are rarely used in standard typesetting but are favored for artistic expression. Regarding font width, there exists a broad spectrum from light to black (which is akin to bold in Latin typography). However, adjusting font weight in CJK characters involves more than merely altering stroke width: The exact boldness of each stroke must be carefully tuned to ensure visual consistency across characters with varying stroke counts, thus requiring more efforts.
For the sake of completeness, we also want to emphasize the regional difference of CJK characters (?). These regional variations in typesetting are handled by both by font design and encoding. Another point of distinction is the use of simplified versus traditional characters, which are represented by different code points. However, these aspects, while important, fall outside the scope of this paper and and we leave it for further study.
@@
Another line of work in generating CJK characters involves transferring from a standardized glyph. Specifically, it has been proposed to generate the glyph in the desired style based on a standard glyph, usually using a reference font that covers CJK code points comprehensively. This approach becomes promising since only a small number of fonts covers a complete set of CJK code points due to significant effort required to produce a consistent appearance for all CJK characters. Examples of such reference fonts that are free to use include Noto CJK (?) and Jigmo (?). The availability of these comprehensive fonts enables the training of a conditional generative model on a limited set of paired glyphs, which can be applied to the entire set of CJK in Unicode. Many of these models are based on Generative Adversarial Networks (GAN) (???????). Some of them leverage additional information about components and composition. While these efforts has seen promising results, they still face quality issues observable upon closer inspection, likely due to the inherent challenges in training GANs.
Perhaps the work that mostly closely related to ours, in that it uses a Denoising Diffusion Probabilistic Model (DDPM) (?) is (?), who introduces Glyph Conditional DDPM (GC-DDPM) that is built on a UNet architecture (?) and generates glyphs from a reference glyph. The key distinction lies in that our method focuses on both typesetting style and brush-based calligraphy generation for font designing and artist, aiming at publisher-level quality suitable for professional font design. In contrast, GC-DDPM aims to synthesize data to enhance the performance of downstream classification on stroke-based hand-writing recognition tasks through data augmentation. Nonetheless, the design decisions made in GC-DDPM have provided valuable insights for our approach.
@@
Style # / Form ${}^{a}$ Font
@@
Regular Script ${}^{b}$ 83534 / P TW-Kai
Serif (Ming/Song) ${}^{b}$ 83534 / P TW-Sung
@@
Serif (Ming/Song) ${}^{c}$ 50217 / P Noto Serif Tangut
Serif (Ming/Song) ${}^{d}$ 22741 / P NomNaTong
${}^{a}$ P for Printed From and C for Calligraphy Form.
${}^{b}$ CNS 11643 Standard. ${}^{c}$ Only for Tangut Script.
${}^{d}$ Contains CJK, including Chu Nom
@@
For training and inference, we use publicly-available fonts since they provide a large set of consistently looking glyphs. Concretely, we use a wide range of free fonts in both printed and calligraphy forms, detailed in Table 1. During training, we construct glyph pairs by looking at the the shared glyphs between the two fonts.
@@
Figure 8: Our method generating CJK characters in a wide range of styles. The upper block is very-commonly-used characters, while the lower block are characters drawn randomly from a collection of less common ones. In each block, we show reference characters in gray, and generated characters in black. Each row shows different styles our method could generate: they are respectively (1) Gothic typefaces, (2) Regular script, (3) Song/Ming in a different standard (CNS 11643), (4) Clerical script, (5-6) two different Semi-cursive scripts and (7) Cursive script. See Table LABEL:fig:tbl:fonts for font data.
@@
To demonstrate the overall generation capability of our proposed method, we present a matrix of generated results in Figure 8, featuring a wide range of fonts from Table 1. Our method applies to both common and rare characters, the latter not present in the training dataset. Specifically, we using Noto Serif TC, a Serif (Ming/Song) font as the reference to generate characters in the desired fonts. These desired fonts include a wide range of styles such as (1) Noto Sans TC (Gothic Typefaces), (2) TW-Kai (Regular Script), (3) TW-Sung (Serif (Ming/Song)), along with calligraphy styles including (4) AoyagiReisho (Clerical Script), (5-6) KouzanMouhitu and KouzanGyousho (Semi-cursive Script), (7) KouzanSousho (Cursive Script). The selection of common characters is based on statistics (?), while less common characters are randomly sampled from Unicode block CJK Unified Ideographs (?), which contains the first batch of $20,992$ CJK characters in Unicode. It is shown that our proposed method works consistently generates high-quality characters across this diverse range of styles, where the results are coherent and recognizable to native speakers.
@@
A great challenge of such scripts is the extremely limited resource due to lack of active users, as Vietnamese nowadays is written in Latin-based alphabet and Tangut language has no living speakers. Two fonts in Tabled 1 represent almost the entirety of free resources available for these scripts, and there is neither demand nor resources to create fonts in other styles for these writing systems.
@@
Figure 14: Inference using different number of steps for diffusion. We here show the results with different number of steps. It could be seen that only the most extreme case of 2 steps we may see a visible artifact. We also compare a wide range of steps by looking at the mean of pixel-wise matching between them and the result from 1000 steps.
@@
The diffusion model is trained and inferred with discretization of time steps from the continuous timespace $[0,1]$ into $T$ steps. Although we train the model with a discretization of $T=1000$ steps, it is suggested (??) that effective inference can be achieved with significantly fewer steps. While we do not explicitly employ techniques to optimize for few-step sampling, exploring the minimum number of steps required for accurate generation remains valuable. As shown in Figure 14, $5$ steps are sufficient and could be treated as feasible approximation. This efficiency allows for generating a character in under one second on a T4 GPU, which is considered less powerful and more suited for inference tasks.

### v2

@@ Abstract
Chinese, Japanese, and Korean (CJK), with a vast number of native speakers, have profound influence on society and culture. The typesetting of CJK languages carries a wide range of requirements due to the complexity of their scripts and unique literary traditions. A critical aspect of this typesetting process is that CJK fonts need to provide a set of consistent-looking glyphs for approximately one hundred thousand characters. However, creating such a font is inherently labor-intensive and expensive, which significantly hampers the development of new CJK fonts for typesetting, historical, aesthetic, or artistic purposes.
@@ Introduction
The importance of East Asian languages that use Chinese characters, including Chinese, Japanese and Korean (CJK) languages, is profound. For example, they are used by more than $1.56$ billion speakers (?) which account for more than $25\%$ of global GDP (?). Also, historically, a wide range of scripts have been heavily influenced by Chinese characters, adding more historical and culture value. In the modern era, with the proliferation of printing and information technology, writing scripts need to address the challenges typesetting: for example, the requirements on the visual appearance of characters means that a font must provide well-designed and consistent-looking glyphs across the glyph repertoire.
However, the unique aspect for CJK typesetting is the sheer number of characters that makes such requirement demande more effort: CJK contains nearly one-hundred thousand unique characters in the latest Unicode 15.1 (?). Furthermore, there is also an emphasis on typefaces categories with different styles [^1] Here “style” is in a general sense rather than its specific meaning in font development as in “style instance”., as they carry practical, cultural, and artistical meaning (???). Consequently, the font industry faces a dilemma: either making one style instance with as wide of coverage as possible, or making fonts with wider style support but with limited coverage. Despite these compromises, font creation remains an laborious and expensive task. Therefore, it remains a critical challenge to efficiently producing fonts that encompass a diverse range of various, multi-style yet legitimate CJK glyphs for the entire set of nearly one hundred thousand characters.
To address these challenges, we propose a diffusion-based method (Figure 1) capable of generating characters in diverse styles conditioned on a single standard reference glyph. The reference glyph only needs to outline the character’s shape, obviating the need for intricate design work. As a diffusion model, it consists of two processes: a diffusion process that gradually adds noise to destroy the data, and a reverse process that generates new samples from scratch by progressively removing noise. A deep neural network is trained to estimate the noise, conditioned on the reference character. Unlike common text-to-image models which must memorize the shapes of characters and thus suffering from limited data of rare characters, our method works uniformly well for all CJK glyphs, from common to rare ones, as long as we have an available reference. Furthermore, by injecting style information in the temporal embedding for diffusion, our method can not only generate characters in various styles but also interpolate between these styles.
@@
The importance of glyphs in typesetting is multifaceted, with style (typefaces) and weight (??) being two crucial aspects. In Latin typesetting, well-recognized styles include the serif (“brown fox”) and sans-serif (“brown fox”), while weight variations are exemplified by bold (“brown fox”) and italics (“brown fox”). CJK typesetting is not an exception, and even encompasses more delicacy regarding this matter: it requires a consistent appearance across a vast number of characters combined with unique cultural practices. Also, CJK typesetting places a significant emphasis on typeface categories, which carry practical, cultural, and artistic significance (???).
Notably, styles in CJK typesetting, while similar, do not entirely align with Latin typesetting practices. They are better shown by several key styles unique to CJK typography in Figure 3: (1) Ming/Song, a printed form style that has evolved from the regular script over centuries, is the most widely used and is akin the serif style in Latin typography. (2) Gothic typefaces, printed form styles that convey a sense of modernity and are akin to sans serif styles in Latin typography. (3) Regular script, a printed form style that mimics handwritten forms, is primarily used for educational purposes and also serves a role similar to italics in Latin typography, although italics per se do not exist in CJK typesetting. (4) Calligraphy form, such as semi-cursive and cursive script, that are rarely used in standard typesetting but are favored for artistic expression. Regarding font width, there exists a broad spectrum from light to black (which is akin to bold in Latin typography). However, adjusting font weight in CJK characters involves more than merely altering stroke width: The exact boldness of each stroke must be carefully tuned to ensure visual consistency across characters with varying stroke counts, thus requiring more effort.
For the sake of completeness, we also want to emphasize the regional difference of CJK characters (?). These regional variations in typesetting are handled by both font design and encoding. Another point of distinction is the use of simplified versus traditional characters, which are represented by different code points. However, these aspects, while important, fall outside the scope of this paper and and we leave them for further study.
@@
Another line of work in generating CJK characters involves transferring from a standardized glyph. Specifically, it has been proposed to generate the glyph in the desired style based on a standard glyph, usually using a reference font that covers CJK code points comprehensively. This approach becomes promising since only a small number of fonts covers a complete set of CJK code points due to significant effort required to produce a consistent appearance for all CJK characters. Examples of such reference fonts that are free to use include Noto CJK (?) and Jigmo (?). The availability of these comprehensive fonts enables the training of a conditional generative model on a limited set of paired glyphs, which can be applied to the entire set of CJK in Unicode. Many of these models are based on Generative Adversarial Networks (GAN) (???????). Some of them leverage additional information about components and composition. While these efforts has seen promising results, they still face quality issues observable upon closer inspection, likely due to the inherent challenges of training GANs.
Perhaps the work that is mostly closely related to ours, in that it uses a Denoising Diffusion Probabilistic Model (DDPM) (?) is (?), which introduces Glyph Conditional DDPM (GC-DDPM) that is built on a UNet architecture (?) and generates glyphs from a reference glyph. The key distinction lies in that our method focuses on both typesetting style and brush-based calligraphy generation for font designing and artist, aiming at publisher-level quality suitable for professional font design. In contrast, GC-DDPM aims to synthesize data to enhance the performance of downstream classification on stroke-based hand-writing recognition tasks through data augmentation. Nonetheless, the design decisions made in GC-DDPM have provided valuable insights for our approach.
@@
Style # / Form a Font
@@
Regular Script b 83534 / P TW-Kai
Serif (Ming/Song) b 83534 / P TW-Sung
@@
Serif (Ming/Song) c 50217 / P Noto Serif Tangut
Serif (Ming/Song) d 22741 / P NomNaTong
a P for Printed From and C for Calligraphy Form.
b CNS 11643 Standard. c Only for Tangut Script.
d Contains CJK, including Chu Nom
@@
For training and inference, we use publicly-available fonts since they provide a large set of glyphs with a consistent look. Concretely, we use a wide range of free fonts in both printed and calligraphy forms, detailed in Table 1. During training, we construct glyph pairs by looking at the the shared glyphs between the two fonts.
@@
Figure 8: Our method generating CJK characters in a wide range of styles. The upper block is very-commonly-used characters, while the lower block are characters drawn randomly from a collection of less common ones. In each block, we show reference characters in gray, and generated characters in black. Each row shows different styles our method could generate: they are respectively (1) Gothic typefaces, (2) Regular script, (3) Song/Ming in a different standard (CNS 11643), (4) Clerical script, (5-6) two different Semi-cursive scripts and (7) Cursive script. See Table 1 for font data.
@@
To demonstrate the overall generation capability of our proposed method, we present a matrix of generated results in Figure 8, featuring a wide range of fonts from Table 1. Our method applies to both common and rare characters, the latter not present in the training dataset. Specifically, we using Noto Serif TC, a Serif (Ming/Song) font as the reference to generate characters in the desired fonts. These desired fonts include a wide range of styles such as (1) Noto Sans TC (Gothic Typefaces), (2) TW-Kai (Regular Script), (3) TW-Sung (Serif (Ming/Song)), along with calligraphy styles including (4) AoyagiReisho (Clerical Script), (5-6) KouzanMouhitu and KouzanGyousho (Semi-cursive Script), (7) KouzanSousho (Cursive Script). The selection of common characters is based on statistics (?), while less common characters are randomly sampled from Unicode block CJK Unified Ideographs (?), which contains the first batch of $20,992$ CJK characters in Unicode. It is shown that our proposed method consistently generates high-quality characters across this diverse range of styles, which are coherent and recognizable to native speakers.
@@
A great challenge of such scripts is the extremely limited resource due to lack of active users, as Vietnamese nowadays is written in Latin-based alphabet and Tangut language has no living speakers. Two fonts in Table 1 represent almost the entirety of free resources available for these scripts, and there is neither demand nor resources to create fonts in other styles for these writing systems.
@@
Figure 14: Inference using different number of steps for diffusion. We here show the results with different number of steps. It could be seen that only the most extreme case of two steps we may see a visible artifact. We also compare a wide range of steps by looking at the mean of pixel-wise matching between them and the result from 1000 steps.
@@
The diffusion model is trained and inferred with discretization of time steps from the continuous timespace $[0,1]$ into $T$ steps. Although we train the model with a discretization of $T=1000$ steps, it is suggested (??) that effective inference can be achieved with significantly fewer steps. While we do not explicitly employ techniques to optimize for few-step sampling, exploring the minimum number of steps required for accurate generation remains valuable. As shown in Figure 14, five steps are sufficient and could be treated as feasible approximation. This efficiency allows for generating a character in under one second on a T4 GPU, which is considered less powerful and more suited for inference tasks.